# Lie symmetry analysis of the two-Higgs-doublet model field equations

### v1

@@ 4.1 A bilinear formalism
The 2HDM potential may also be written in a particularly useful manner by gauge invariant bilinears [51, 52].
@@
where $a\in\{1,2,3\}$ and bilinears $r_{a}$ are given in terms of the Pauli matrices $\sigma_{a}$, which infer
@@
To keep the potential invariant, the coupling constants then transforms as follows under this change of basis,
@@
Now, since $\text{Ad}_{\mathsf{SU}(2)}=\mathsf{SO}(3)$ all matrices of $\mathsf{SO}(3)$ correspond to a $\mathsf{SU}(2)$ Higgs basis transformation, and we may always diagonalize the matrix $\Lambda$ of (4.17) by a basis transformation. Given a arbitrary potential, we may change basis in such a way that $\Lambda$ is diagonalized. This means that
@@
in the new basis. The basis shift will of course depend on the values of the parameters of $\Lambda$ in the original basis, and we have hence reduced the number of parameters by three. A slightly different bilinear formalism was applied in the beautiful derivation of the strict variational symmetries of the 2HDM in [21]. A crucial part of the proof is the aforementioned diagonalizability of $\Lambda$, which does not carry over to the analogous matrix $\Lambda$ for higher $N$, since $\text{Ad}_{\mathsf{SU}(N)}\varsubsetneq\mathsf{SO}(N^{2}-1)$ is dramatically smaller when $N>2$. This again makes classification of symmetries by bilinear formalisms much harder for $N>2$.
@@ 4.2 Finding and solving determining equations
We now find the determining equations (2.17) of the system of PDEs given by (4.6), by applying SYM [45]. The system (4.6) yields a set of 1733 determining equations, all linear PDEs in the coefficients $\xi^{\mu}$ and $\eta^{i}$ of (2.6). Since we are only interested in scalar symmetries, we set
@@
$\eta^{i}$ $=0,\quad\text{for all}\quad 9\leq i\leq 24,$ (4.25)
which means spacetime variables and gauge bosons are kept constant under the considered transformations. Then the simplest (shortest) of the remaining determining equations read
@@
which means the $\eta^{i}$’s are affine (or sometimes referred to as linear) in the scalar fields, and do not depend on the gauge fields. Hence, we may proceed solving the determining equations by substituting
@@
with an implicit sum over $j$ from 1 to 8, into the determining equations. We don’t include explicit spacetime variables in (4.28), since we are only considering pure scalar symmetries, although it would have been a principal possibility for non-scalar symmetries. After substituting (4.28), we are left with a system of 353 polynomial equations
$P_{k}(y^{1},\ldots,y^{q})=0$ (4.29)
in the $q=24$ fields, with $1\leq k\leq 353$. The polynomials $P_{k}$ will contain parameters from the Lagrangian, in addition to the $a_{i}$’s and $b_{ij}$’s. All the equations (4.29) have to hold for all values of the fields, to have a symmetry. This means that the coefficients of all distinct monomials in the fields, in all equations (4.29), have to be zero. More precisely, let a general monomial $\bar{m}$ in the $q=24$ fields be written
@@
when $C\equiv C_{n_{1},\ldots,n_{q}}(m^{2},\lambda,g,g^{\prime}Y)$ is the maximal coefficient of $\bar{m}(n_{1},\ldots,n_{q})$, that is
@@
for any, non-zero coefficient $D\equiv D_{n_{1},\ldots,n_{q}}(m^{2},\lambda,g,g^{\prime}Y)$. This splitting in monomials yields 1412 equations of the form $C=0$, some of them very simple. The 80 simplest equations are of the type
@@
parameters which do not vanish at this stage. We also could have considered the case $g^{\prime}=0$ here, and hence investigated custodial symmetries [53, 54, 34, 32, 55, 56, 57, 58], but we will refrain from it, as these of course are not exact symmetries of the considered Lagrangian (4.1), and as this could have doubled up the analysis. Eq. (4.34) means there cannot be any pure scalar shift symmetries of the field equations in a 2HDM. We continue our process of solving the determining equations by substituting the solutions (4.34) and (4.35) into the 1412 determining equations, where only the parameters (4.2) are kept non-zero among the $b_{ij}$’s.
At this stage we also implement (4.5), i.e. set $m_{12}^{2}=0$ and $\Im(\lambda_{5})=0$, although we just as well could have done it from the start, before calculating the Euler-Lagrange equations. We then end up with a system
@@ 4.3 Parameter cases and reductions of the potential
Let a reduced potential be a potential where one or more parameters have been eliminated, by transforming the potential to a specific scalar basis, through an $\mathsf{U}(2)$ Higgs basis transformation (4.4). Reduced potentials have been discussed and employed in e.g. [13, 59, 21, 32]. We proceed by considering different parameter cases, and will try to keep the potential optimally reduced in each case. This means that we eliminate as many parameters as possible through an appropriate basis transformation, and hence, more or less, lock the basis and thereby discard reparametrization freedom. We will do this to avoid that the same symmetry may manifest itself in different bases of the doublets excessively, typically by Lie algebras with generators involving potential parameters. Although all these manifestations of the same symmetry will be detected by the Lie symmetry analysis (for a clarification of this, see Appendix B), we want to avoid this as much as possible, because we otherwise would have to show that two different manifestations of the same symmetry in fact are equivalent, which may imply extra work.
@@ 4.3.1 The case $m_{11}^{2}\neq m_{22}^{2}$
In addition to the diagonalization of the mass-squared matrix, and hence the reduction (4.5) of the potential, we now assume
@@
for $h>0$ cf. (4.33) and certain indices $i,j,k$ and $l$. Solving these equations gives
@@
which means we have only two free parameters $b_{21}$ and $b_{78}$. Inserting (4.3.1) into the full set of determining equations, and then applying Mathematica’s built-in Reduce function, gives us that either
@@
Here, the case (4.42) substituted into (4.28) and then into the infinitesimal generator (2.6) results in a 1-dimensional algebra that is present for all [^10] This corresponds to the fact there are no restricitions on the parameters of the potential in (4.42). potentials, given by the generator
@@
which again infers the symmetry algebra
@@
the Lie algebra of the hypercharge symmetry group $\mathsf{U}(1)_{Y}$. The only other possibility (4.43) corresponds to a 2-dimensional algebra, parametrized by the two now free parameters $b_{21},b_{78}$, since there are no restrictions on $b_{21},b_{78}$ in (4.43). Eqs. (4.28) and (2.6) then gives a symmetry algebra
@@
which spans the Lie algebra $\mathfrak{u}(1)_{\text{PQ}}$ of the Peccei-Quinn $\mathsf{U}(1)$ symmetry. We may demonstrate this is a strict variational symmetry by applying the fact $\operatorname{pr}X_{\text{PQ}}(T)=0$ (i.e. the Peccei-Quinn symmetry is a SVS of the kinetic terms), together with Theorem 1. Alternatively show (2.21) holds for potential parameters (4.5) and (4.43), or just check the consistency with the results in [21] or e.g. Table 5 of [26].
@@ 4.3.2 The case $m_{11}^{2}=m_{22}^{2}$
and substitute (4.48) in the determining equations (4.37). The 148 simplest of these equations are of the form (4.40) again, and we solve and substitute the resulting equations in the total system of determining equations. The only free parameters $b_{ij}$ in the determining equations can now be taken as
@@
Because of (4.48), the mass-squared matrix will remain diagonal under any additional Higgs basis transformations, hence we have freedom to reduce the potential some more, without introducing a new $m_{12}^{2}$ parameter. We therefore choose to diagonalize the matrix $\Lambda$ of (4.17), which means
@@
since $\Re(\lambda_{5})=0$ would also have let us reduce the potential by eliminating $\Im(\lambda_{6})$. By substituting (4.50) into the determining equations and solving for the parameters (4.49), with the assumption (4.51), we get 3 different solutions for the parameters $B_{2}$, cf. (4.49), and hence 3 possible Lie symmetry algebras: Let
@@
Then the Lie algebra of the reparametrization group $\mathsf{SU}(2)$ (i.e. the basis transformations), denoted $\mathfrak{su}(2)_{\text{HF}}$ (where ”HF” stands for ”Higgs Family”), will be spanned by the latter generators and
@@
with $b_{51}$ free, and this holds for the displayed parameter conditions, of course in addition to the other conditions like (4.5). The two other solutions are
@@
where both solutions hold for several, different conditions on the potential parameters, similar to the ones given in (4.3.2). Then the 3 symmetry algebras are respectively
@@
We can conclude all these algebras are strict variational symmetries, without considering all the distinct sets of conditions on the parameters of the potential: The prolonged infinitesimal elements $\operatorname{pr}X$ of all generators $X$ in (4.58) will annihilate the kinetic terms when applied on the most general $\mathcal{L}$. This may be checked explicitly, but it also follows from the well-known fact that $\mathsf{SU}(2)_{\text{HF}}$ is a strict variational symmetry group of the kinetic terms of the 2HDM, see [34] for a proof that the strict variational symmetries of the kinetic terms are $\mathfrak{su}(N)$ for a general NHDM. Then, Theorem 1 guarantees these symmetries are strictly variational, independent of the conditions on the parameters of the potential. Moreover, we will in section 4.3.3 show that the symmetries given by $\mathfrak{g}_{\text{I}}$ and $\mathfrak{g}_{\text{II}}$ are equivalent, which means that there is a new Higgs basis where $\mathfrak{g}_{\text{I}}$ will be the same as $\mathfrak{g}_{\text{II}}$ was in the old basis.
@@
in addition to (4.48) and (4.50). By a rephasing of the doublets, we may now eliminate $\Im(\lambda_{6})$,
@@
without introducing new parameters.
@@
If $\lambda_{1}+\lambda_{2}=2(\lambda_{3}+\lambda_{4})$ the matrix $\Lambda$ will be proportional to identity, and we can perform $\mathsf{SO}(3)$ rotations without altering the elements of $\Lambda$. Then we may freely rotate $L$ while $M$ remains zero, such that $L$ points in a desired direction, e.g. in the $x$-direction, and then will $\lambda_{1}=\lambda_{2}$. Hence, we will first consider the case where
@@
in addition to all the previous assumptions, like (4.61) and (4.62). By solving the determining equations for this case, we get 3 solutions for the surviving parameters $B_{2}$, cf. (4.49). In the same manner as in earlier sections, the resulting Lie symmetry algebras are
@@
where two different solutions for the $b$’s correspond to the same algebra $\mathfrak{u}(1)_{Y}$, while $\mathfrak{g}_{1}$ is isomorphic to to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. In the same manner as before, all symmetries given by (4.64) are, by Theorem 1, strictly variational, since they annihilate the kinetic terms of the 2HDM.
@@
We now consider the last case with, as explained in the beginning of the last paragraph,
@@
in addition to (4.48), (4.50), (4.61) and (4.62). Under these conditions we get two solutions for the $b$’s. In the first solution, $\Re(\lambda_{6})$ is free, and the corresponding symmetry algebra is
@@
which again is isomorphic to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. The second solution is that all the four parameters of $B_{2}$ are free, cf. (4.49), while $\Re(\lambda_{6})=0$. In this case the symmetry algebra is

### v2

@@ 4.1 A bilinear formalism
The 2HDM potential may also be written in a particularly useful manner by gauge invariant bilinears, a formalism first introduced in [50] and subsequently developed for the general 2HDM in [51].
@@
where $a\in\{1,2,3\}$ and bilinears $r_{a}$ are given in terms of the Pauli matrices $\sigma_{a}$, which imply
@@
To keep the potential invariant, the coupling constants are transformed as follows under this change of basis:
@@
Because $\text{Ad}_{\mathsf{SU}(2)}=\mathsf{SO}(3)$ all matrices of $\mathsf{SO}(3)$ correspond to an $\mathsf{SU}(2)$ Higgs basis transformation, and we may always diagonalize the matrix $\Lambda$ of (4.17) by a basis transformation. Given an arbitrary potential, we can change the basis in such a way that $\Lambda$ is diagonalized. This means that
@@
in the new basis. The basis shift will, of course, depend on the values of the parameters of $\Lambda$ in the original basis, hence we have reduced the number of parameters by three. A slightly different bilinear formalism was applied in the beautiful derivation of the strict variational symmetries of the 2HDM in [21]. A crucial part of the proof is the aforementioned diagonalizability of $\Lambda$, which does not carry over to the analogous matrix $\Lambda$ for higher $N$, because $\text{Ad}_{\mathsf{SU}(N)}\varsubsetneq\mathsf{SO}(N^{2}-1)$ is dramatically smaller when $N>2$. This again makes the classification of symmetries using bilinear formalisms much harder for $N>2$.
@@ 4.2 Finding and solving determining equations
We now find the determining equations (2.19) of the system of PDEs given by (4.6) by applying SYM [43]. This yields a system of 1733 determining equations for the coefficients $\xi^{\mu}$ and $\eta^{i}$ of the point-symmetry generator (2.8).
These determining equations follow from the symmetry condition (2.19), that is, by computing $\operatorname{pr}X\big(E_{i}(\mathcal{L})\big)$ for all $i$ as functions on the relevant jet space and then imposing $E(\mathcal{L})=0$ by substituting 24 of the highest-order derivatives, cf. (3.4). In geometric terms, this ensures that the change of the functions $E_{i}(\mathcal{L})$ along the direction $\operatorname{pr}X$ vanishes when starting on the solution manifold $\mathcal{M}_{E(\mathcal{L})}$, so that $\operatorname{pr}X$ is tangent to $\mathcal{M}_{E(\mathcal{L})}$. Hence $X$ generates a Lie point symmetry, since it moves us infinitesimally from one solution to another. Finally, by demanding that the coefficients of all linearly independent monomials in the remaining derivatives of the fields vanish, we obtain a linear overdetermined system of PDEs for the functions $\xi^{\mu}$ and $\eta^{i}$, cf. (3.1).
As we are interested only in scalar symmetries, we set
@@
$\eta^{i}$ $=0,\quad\text{for all}\quad 9\leq i\leq 24.$ (4.25)
This means that the spacetime variables and gauge bosons are kept constant under the considered transformations. Then the simplest (shortest) of the remaining determining equations are
@@
which means that the $\eta^{i}$’s are affine (or sometimes referred to as linear) in the scalar fields and do not depend on the gauge fields. For strict variational symmetries, (4.26) is natural: any quadratic or higher dependence $\eta^{i}=\mathcal{O}(\phi^{2})$ would generate higher-order monomials (such as $\phi^{5}$ or $\phi^{8}$) in the transformed potential, which would spoil the invariance of the potential.
Hence, we can proceed by solving the determining equations by substituting
@@
with an implicit sum over $j$ ranging from 1 to 8 into the determining equations. We do not include explicit spacetime variables in (4.28) because we are only considering pure scalar symmetries, although it would have been a principal possibility for non-scalar symmetries. After substituting (4.28), we obtain a system of 353 polynomial equations
$P_{k}(y^{1},\ldots,y^{q})=0,$ (4.29)
for the $q=24$ fields (4.7), with $1\leq k\leq 353$. All derivatives of the fields were eliminated when deriving the general determining equations for the functions $\eta^{i}$ and $\xi^{\mu}$. The polynomials $P_{k}$ depend on the parameters of the Lagrangian, as well as on the unknown constants $a_{i}$ and $b_{ij}$, which we wish to determine for the different possible choices of Lagrangian parameters. For a symmetry to be present, all equations (4.29) must hold for all field values, in complete analogy with the requirement that (3.1) must hold for all values of $\phi$ (and $x$). This implies that, in each equation (4.29), the coefficients of all distinct monomials in the fields must vanish.
More precisely, let a general monomial $\bar{m}$ in the $q=24$ fields be written as:
@@
when $C\equiv C_{n_{1},\ldots,n_{q}}(a,b,m^{2},\lambda,g,g^{\prime}Y)$ is the maximal coefficient of $\bar{m}(n_{1},\ldots,n_{q})$, that is
@@
for any, non-zero coefficient $D\equiv D_{n_{1},\ldots,n_{q}}(a,b,m^{2},\lambda,g,g^{\prime}Y)$. This splitting into monomials yields 1412 equations of the form $C=0$, some of which are very simple. The 80 simplest equations are of this type:
@@
parameters that did not vanish at this stage. We also could have considered the case $g^{\prime}=0$ here, and hence investigated custodial symmetries [52], but we will refrain from it, as these are not exact symmetries of the considered Lagrangian (4.1), and as this could have doubled the analysis. [^10] See [53] for conditions for the canonical custodial symmetry (CCS) in the 2HDM, while [32] classifies non-canonical 2HDM symmetries. The Lie symmetry analysis used in this work would detect all such symmetries, including the non-canonical ones.
Equation (4.34) means that there cannot be any pure scalar shift symmetries of the field equations in a 2HDM. We continue our process of solving the determining equations by substituting the solutions (4.34) and (4.35) into the 1412 determining equations, where only the parameters (4.2) are kept non-zero among the $b_{ij}$’s.
At this stage, we also implement (4.5), that is, set $m_{12}^{2}=0$ and $\Im(\lambda_{5})=0$, although we just as well could have done so from the start, before calculating the Euler-Lagrange equations. Then, we end up with a system
@@ 4.3 Parameter cases and reductions of the potential
Let a reduced potential be a potential in which one or more parameters have been eliminated by transforming to a specific scalar basis through a $\mathsf{U}(2)$ Higgs-basis transformation (4.4). Such reduced potentials have been used, for example, in early studies of CP violation in the 2HDM [15], in basis-independent analyses [54], in the Minkowski-space formulation of the Higgs potential [21], and in tensor-product-based classifications of scalar symmetries [32].
We proceed by considering different parameter cases, and try to optimally reduce the potential in each case. This means that we eliminate as many parameters as possible through an appropriate basis transformation, and hence, more or less, lock the basis and thereby discard reparametrization freedom. We do this to avoid that the same symmetry manifests itself in different bases of the doublets excessively, typically by Lie algebras with generators involving potential parameters. Although all these manifestations of the same symmetry will be detected by the Lie symmetry analysis (for a clarification of this, see Appendix B), we want to avoid this as much as possible, because we would otherwise have to show that two different manifestations of the same symmetry are equivalent, which may imply extra work.
@@ 4.3.1 The case $m_{11}^{2}\neq m_{22}^{2}$
In addition to the diagonalization of the mass-squared matrix, and hence the reduction (4.5) of the potential, we now assume that
@@
for $h>0$ cf. (4.33) and certain indices $i,j,k$ and $l$. Solving these equations yields:
@@
which means that we have only two free parameters: $b_{21}$ and $b_{78}$. Inserting (4.3.1) into the full set of determining equations and then applying Mathematica’s built-in Reduce function, we obtain
@@
Here, substituting (4.42) into (4.28) and then into the infinitesimal generator (2.8) results in a 1-dimensional algebra that is present for all [^11] This corresponds to the fact that there are no restrictions on the parameters of the potential in (4.42). potentials, given by the generator
@@
which again implies the symmetry algebra
@@
that is, the Lie algebra of the hypercharge symmetry group $\mathsf{U}(1)_{Y}$. The only other possibility (4.43) corresponds to a 2-dimensional algebra, parametrized by the two now free parameters $b_{21},b_{78}$, because there are no restrictions on $b_{21},b_{78}$ in (4.43). Equations (4.28) and (2.8) then yield a symmetry algebra
@@
which spans the Lie algebra $\mathfrak{u}(1)_{\text{PQ}}$ of the Peccei-Quinn $\mathsf{U}(1)$ symmetry. We demonstrate that this is a strict variational symmetry by applying the fact $\operatorname{pr}X_{\text{PQ}}(T)=0$ (i.e., the Peccei-Quinn symmetry is an SVS of the kinetic terms), together with Theorem 1. Alternatively, we can show (2.23) holds for potential parameters (4.5) and (4.43), or we can simply check the consistency with the results in [21] or Table 5 of [26].
@@ 4.3.2 The case $m_{11}^{2}=m_{22}^{2}$
and substituting (4.48) in the determining equations (4.37). The 148 simplest of these equations are of the form (4.40), and we solve and substitute the resulting equations in the total system of determining equations. The only free parameters $b_{ij}$ in the determining equations can now be taken as:
@@
Because of (4.48), the mass-squared matrix will remain diagonal under any additional Higgs basis transformations, hence, we have the freedom to further reduce the potential, without introducing a new $m_{12}^{2}$ parameter. Therefore, we choose to diagonalize the matrix $\Lambda$ in (4.17), which means that
@@
since $\Re(\lambda_{5})=0$ would also have let us reduce the potential by eliminating $\Im(\lambda_{6})$. By substituting (4.50) into the determining equations and solving for the parameters (4.49), with assumption (4.51), we obtain three different solutions for the parameters $B_{2}$, cf. (4.49), hence three possible Lie symmetry algebras: Let
@@
Then, the Lie algebra of the reparametrization group $\mathsf{SU}(2)$ (i.e. the basis transformations), denoted $\mathfrak{su}(2)_{\text{HF}}$ (where ”HF” stands for ”Higgs Family”), will be spanned by the latter generators and
@@
with $b_{51}$ free, and this holds for the displayed parameter conditions, in addition to the other conditions as in (4.5). The other two solutions are as follows:
@@
where both solutions hold for several, different conditions on the potential parameters, similar to those given in (4.3.2). Then, the three symmetry algebras are as follows:
@@
We can conclude that all these algebras are strict variational symmetries, without considering all the distinct sets of conditions on the parameters of the potential: The prolonged infinitesimal elements $\operatorname{pr}X$ of all generators $X$ in (4.58) will annihilate the kinetic terms when applied to the most general $\mathcal{L}$. This may be verified explicitly, but it also follows from the well-known fact that $\mathsf{SU}(2)_{\text{HF}}$ is a strict variational symmetry group of the kinetic terms of the 2HDM, see [34] for a proof that the strict variational symmetries of the kinetic terms are $\mathfrak{su}(N)$ for a general NHDM. Then, Theorem 1 guarantees that these symmetries are strictly variational, independent of the parameter constraints of the potentials. Moreover, in Section 4.3.3, we will show that the symmetries given by $\mathfrak{g}_{\text{I}}$ and $\mathfrak{g}_{\text{II}}$ are equivalent, which means that there is a new Higgs basis where $\mathfrak{g}_{\text{I}}$ will be the same as $\mathfrak{g}_{\text{II}}$ in the old basis.
@@
in addition to the conditions (4.48) and (4.50). By a rephasing of the doublets, we may now eliminate $\Im(\lambda_{6})$,
@@
without introducing any new parameters.
@@
If $\lambda_{1}+\lambda_{2}=2(\lambda_{3}+\lambda_{4})$ the matrix $\Lambda$ will be proportional to identity, and we can perform $\mathsf{SO}(3)$ rotations without altering the elements of $\Lambda$. Then, we can freely rotate $L$ while $M$ remains zero, such that $L$ points in a desired direction, for example, in the $x$-direction, and then will $\lambda_{1}=\lambda_{2}$. Hence, we first consider the case in which:
@@
in addition to all the previous assumptions, such as (4.61) and (4.62). By solving the determining equations for this case, we obtain three solutions for the surviving parameters $B_{2}$, cf. (4.49). In the same manner as in earlier sections, the resulting Lie symmetry algebras are
@@
where two different solutions for the $b$’s correspond to the same algebra $\mathfrak{u}(1)_{Y}$, while $\mathfrak{g}_{1}$ is isomorphic to to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. In the same manner as before, all symmetries given by (4.64) are, by Theorem 1, strictly variational because they annihilate the kinetic terms of the 2HDM.
@@
We now consider the last case with, as explained at the beginning of the last paragraph,
@@
in addition to (4.48), (4.50), (4.61) and (4.62). Under these conditions we obtain two solutions for the $b$’s. In the first solution, $\Re(\lambda_{6})$ is free, and the corresponding symmetry algebra is
@@
which again is isomorphic to $\mathfrak{u}(1)\oplus\mathfrak{u}(1)_{Y}$. The second solution is that all the four parameters of $B_{2}$ are free, cf. (4.49), while $\Re(\lambda_{6})=0$. In this case, the symmetry algebra is